I'm having difficulty with this problem.
-6(5X+2)=198 I was wondering if you could solve this for me and show your work so that I can learn how to do it. I am working with my 8th grade daughter and I wish I could show her how to do this. Maybe if I see how it is done I will be able to explain it to her. I've been working on this with her for an hour. Please Help! Thanks!
Karen

Answer by checkley71(8403)

-6(5X+2)=198
-30X-12=198
-30X=198+12
-30X=210
X=210/-30
X=-7
PROOF
-6(5*-7+2)==198
-6(-35+2)=198
-6(-33)=198
198=198